Circular No. 6818

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

XTE J0055–724 = 1SAX J0054.9–7226 F. E. Marshall, Goddard Space Flight Center (GSFC); and J. C.

Lochner, GSFC and Universities Space Research Association (USRA), report the discovery of a transient x-ray source, designated XTE J0055–724, that was pulsating with a period of 59.0 ± 0.2 s during an observation of the Small Magellanic Cloud (SMC) with the RXTE on Jan. 20: “Based on scans across the source on Jan. 22, we estimate a position of α = 0^{h}55^{m}.3, δ = −72^{o}29′ (equinox 2000.0; estimated error circle 10′). The average flux was ∼ 6 × 10^{−11} erg s^{−1} cm^{−2} in the band 2–10 keV, with approximately sinusoidal variations of ± 30 percent. The source was not seen during an observation on Jan. 12.”

A. Santangelo and G. Cusumano, Istituto di Fisica Cosmica ed Applicazioni dell’Informatica del CNR, Palermo; G. L. Israel, Osservatorio Astronomico di Roma; D. Dal Fiume, M. Orlandini, and F. Frontera, Istituto di Tecnologie e Studio delle Radiazioni Extraterrestri, Bologna; A. N. Parmar, Space Science Department, European Space Agency, ESTEC; F. E. Marshall, Goddard Space Flight Center (GSFC); and J. C. Lochner and R. H. D. Corbet, GSFC and Universities Space Research Association, report: “A BeppoSAX Narrow Field Instruments Target Of Opportunity observation in the direction of the SMC was performed on Jan. 28. The BeppoSAX MECS detected a pulsating source located at α = 0^{h}54^{m}54^{s}.8, δ = −72^{o}26′41′′ (equinox 2000.0; uncertainty radius 50′′), which we have designated as 1SAX J0054.9–7226. The source had a flux of 1.9 × 10^{−11} erg s^{−1} cm^{−2} (2–10 keV) and exhibited pulsations of period 58.969 ± 0.001 s. Our position for 1SAX J0054.9–7226 is consistent with that of the ROSAT WGACAT source 1WGA J0054.9–7226 and of the Einstein source 2E 0053.2–7242. A periodicity search of archival data from the October 1991 ROSAT observation of 1WGA J0054.9–7226 reveals a strong pulsation with period 59.072 ± 0.001 s; the same source was not detected in four ROSAT PSPC pointings performed between April and November 1993. Our preliminary conclusion is that 2E 0053.2–7242 = XTE J0055– 724 = 1WGA J0054.9–7226 is a newly-recognized 59-s transient x-ray pulsar that is likely to be in a Be/high-mass-x-ray-binary system and unrelated to the three x-ray pulsars reported on IAUC 6788, 6803, and 6814. At least four other sources are now in the MECS field-of-view. Analysis of these sources is currently underway.”

1998 February 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green